Shift Summary, Jan 23, evening shift


Subject: Shift Summary, Jan 23, evening shift
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:38:36 CST


       Shift Summary for walk-thru evening shift, 01/23/2000

On Walk-Thru 'Shift': D. Jensen

Hours of beam available at all intensities: 0.0
Hours of beam available at nominal intensity: 0.0
Hours of E799 data (special runs not included): 0.0
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time:
     Beam off - convert to 20 marad targeting angle

========================================================================
Runs ended this shift: none

Tapes brought to FCC this shift:
  (including tapes for online split)

none
========================================================================
Accesses during shift: open

========================================================================
Overview of shift:
 4 pm - all is quiet. Shanahan doing a tour...

 7 pm - Found Libert in high temperature alarm. Chilled water not
         going. Duty Eng came and reset it. Will be watching it for
         a while. Temps on Lieberts up to 85 - 90 deg F. Shut down
         DP for a while.
 8 pm Cool down of the building is working. Vacuum system OK.
         Gas check done. OK
           
 10 pm All is cool this time. Chiller continues to chill. Amazing
         how when it is 7 deg F outside (now down to 4.7), everything
         can go down the tubes because a chiller turns off!
         
 12 pm All remains cool. Outside down to 3.5 degF.

........................................................................
For reference: annunciator pannel on Sourth wall - known problems:
 lights (1,6), (2,6), (6,8) on
        (5,5) blinking ( cold gas room )

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
no running

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

========================================================================
Etc...
  Once again the lack of a copy machine is noted..........
=======================================================================



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:38:39 CST